# SMReferences for Appendix

### v3

@@ 1 Introduction
Before moving forwards, we outline some notation that is used throughout. We denote $\theta$ as the nuisance parameters. All expectations and probabilities (e.g. $\mathbb{E}$ and $\mathbb{P}$) are understood to be under the true data generating distribution unless stated otherwise (in that case, we will attach $\theta$ in the subscripts (e.g. $\mathbb{E}_{\theta}$ and $\mathbb{P}_{\theta}$) to emphasize the dependence of the distribution on $\theta$). As will be clear later in Section 2, we may split the entire dataset into three parts, a main dataset $\mathcal{D}$ of size $n$, an auxiliary dataset $\mathcal{D}_{\mathrm{aux}}$ of size $n_{\mathrm{aux}}$, and a training dataset $\mathcal{D}_{\mathrm{tr}}$ of size $n_{\mathrm{tr}}$. Bold letters will be preserved for to observed data or related quantities over a given dataset. For instance, $\mathbf{T}\equiv(T_{1},\cdots,T_{n})^{\top}$ and $\mathbf{X}=(X_{1}^{\top},\cdots,X_{n}^{\top})^{\top}$, respectively, denote the treatment variables and the covariates over the main dataset. The quantities computed from a given dataset will be attached, in the subscripts, by the index of that dataset. For example, the $n_{\mathrm{tr}}\times p$ covariate matrix from the training dataset will be denoted as $\mathbf{X}_{\mathrm{tr}}$, and similarly for $\mathbf{X}_{\mathrm{aux}}$. Finally, given two vectors $v_{1},v_{2}$ of the same length, $v_{1}\odot v_{2}$ denotes the component-wise multiplication between $v_{1}$ and $v_{2}$.
@@
The rest of this paper is organized as follows. In Section 2, we illustrate our new methodology for ATE, with the theoretical analysis detailed in Section 3. Section 4 extends our new approach to (1) the regression coefficient estimation in a high-dimensional partially linear model and (2) the approximately sparse GLMs considered in Belloni et al. (2014); Chernozhukov et al. (2018); Bradic et al. (2019a) and Smucler et al. (2019). Lastly, Section 5 concludes this article and discusses some future directions. Some of the technical details are deferred to the Appendix.
@@
Chernozhukov et al. (2022) $s_{\pi}\wedge s_{r}=o(\sqrt{n}/\log p)$
@@ 2 Average treatment effects: The Double-Calibration estimator
In this section, we illustrate our new “Double-Calibration (DCal)” methodology for ATE estimation. We observe data $\{O_{i}\coloneqq(X_{i},T_{i},Y_{i})\}_{i=1}^{N}$ that are i.i.d. draws from some underlying data-generating distribution $\mathbb{P}_{\theta^{\ast}}$ where from this section forward $\theta^{\ast}=(\pi^{\ast},r^{\ast})$ denotes the true OR and PS models. We assume throughout that, as standard in the literature on treatment effect estimation, $\mathbb{P}_{\theta^{\ast}}$ satisfies the unconfoundedness assumption:
@@
and the overlap/positivity assumption:
@@
It is well-known that, under Consistency, Conditions 1 and 2 (Hernán and Robins, 2023), the parameter of interest, ATE, can be identified as $\tau^{\ast}\equiv\mathbb{E}[r^{\ast}(X)]\equiv\mathbb{E}\left[\frac{TY}{\pi^{\ast}(X)}\right]\equiv\mathbb{E}[H(\theta^{\ast})]$. We further impose the following modeling assumption on the nuisance parameters $\theta^{\ast}=(\pi^{\ast},r^{\ast})$:
@@ 2.1 Preliminary: Singly-calibrating the OR model
$\hat{\tau}_{\textrm{SCal},r}\coloneqq\frac{1}{n}\sum_{i=1}^{n}H_{i}(\hat{\pi},\tilde{r})\equiv\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\tilde{r}_{i})}{\hat{\pi}_{i}}+\tilde{r}_{i}\right),\text{ where }\tilde{r}_{i}\coloneqq\hat{r}_{i}+\hat{\mu}_{i}.$
@@
$\hat{\bm{\mu}}=$ $\;{\arg\min}_{\bm{\mu}\in\mathbb{R}^{n}}\frac{1}{n}\|\bm{\mu}\|_{2}^{2}$ (4)
$\mathrm{s.t.}\;\;$ $\left\|\frac{1}{n_{\mathrm{aux}}}{\mathbf{X}}_{\mathrm{aux}}^{\top}{\bm{\Pi}}_{\mathrm{aux}}\tilde{\mathbf{R}}_{\mathrm{aux}}-\frac{1}{n}{\mathbf{X}}^{\top}{\bm{\Pi}}{\bm{\mu}}\right\|_{\infty}\leq\eta_{r},$ (5)
$\left\|{\bm{\mu}}\right\|_{\infty}\leq M_{r}$ (6)
where $M_{r}>0$ is some known constant that depends on the bound on $\hat{r}$ and $r^{\ast}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the covariate matrix, ${\bm{\Pi}}\in\mathbb{R}^{n\times n}$ is a diagonal matrix whose $(i,i)$-th entry is $\phi^{\prime}(X_{i}^{\top}\hat{\gamma})/\hat{\pi}_{i}$ and $\tilde{\mathbf{R}}$ is the $n$-dimensional residual vector whose $i$-th element is $\tilde{R}_{i}\coloneqq T_{i}/\hat{\pi}_{i}\cdot(Y_{i}-\hat{r}_{i})$. Recalling from the Notation section, the quantities $\mathbf{X}_{\mathrm{aux}}$ etc. are defined analogously on the auxiliary dataset $\mathcal{D}_{\mathrm{aux}}$.
@@
$\hat{\tau}_{\textrm{SCal},r}$ is the same as the DML estimator $\hat{\tau}_{\textrm{DML}}$ except $\hat{r}$ is replaced by the post-calibration OR estimator $\tilde{r}=\hat{r}+\hat{\mu}$. It is in fact a variant of the recently proposed DIPW estimator for ATE (Wang and Shah, 2020). Constraint (5) balances $\tilde{r}$ such that it is on average close to $Y$, gauged by covariates $X$ under certain data-dependent weights. The use of the auxiliary dataset $\mathcal{D}_{\mathrm{aux}}$ is to preserve the following moment equation satisfied by the calibrated OR estimator $\tilde{r}$:
@@
because $\hat{\mu}$’s, as constructed, only depends on $X$, but not on $T$ or $Y$, from the main dataset. Note that (7) in fact holds with any quantity, that is independent of $T_{i}$ given $X_{i}$ from the main dataset $\mathcal{D}$, in place of $\hat{\mu}_{i}$. The above moment equation plays a key role in the analysis of the statistical property of $\hat{\tau}_{\textrm{SCal},r}$ in Wang and Shah (2020). ∎
Using the analysis in the proof of Wang and Shah (2020, Theorem 2), since $\hat{\beta}$ and $\hat{\gamma}$, computed using $\mathcal{D}_{\mathrm{tr}}$, are independent from the main sample $\mathcal{D}$, by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}}$, $\hat{\tau}_{\textrm{SCal},r}$ can be shown to be a $\sqrt{n}$-consistent estimator of $\bar{\tau}^{\ast}$ whenever the PS-sparsity $s_{\pi}\ll\sqrt{n}/\log p$, under Conditions 1 to 3.
@@ 2.2 Double calibration
$\hat{\tau}_{\textrm{DCal}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}H_{i}(\tilde{\pi},\tilde{r})\equiv\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\tilde{r}_{i})}{\tilde{\pi}_{i}}+\tilde{r}_{i}\right).$
@@
$\tilde{\gamma}\coloneqq$ $\ {\arg\min}_{\gamma}\|\gamma-\hat{\gamma}\|_{1}$ (8)
@@
$\left|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}}-1\right)\hat{\mu}_{i}\right|\leq\eta_{\pi_{1}}\cdot\frac{\|\hat{\bm{\mu}}\|_{2}}{\sqrt{n}}$ (10)
@@
$\max_{i}T_{i}/\pi_{i}\leq M_{\pi}.$ (12)
In this program, $\pi_{i}\equiv\pi(X_{i})\coloneqq\phi(\tilde{X}_{i}^{\top}\gamma)$ and $M_{\pi}>0$ is some known constant that depends on $c_{\pi}$ [^7] In (9), we use the convention that a function $g:\mathbb{R}^{p}\mapsto\mathbb{R}$ applied to a matrix ${\bm{M}}\in\mathbb{R}^{n\times p}$ ($p$ could be equal to $1$) is a $n$-dimensional vector formed by applying $g$ to each row of the matrix..
Without any assumptions on the sparsity of the PS or the OR model, (8) is always guaranteed to have a feasible solution:
@@
If Condition 2 holds and almost surely, $\mathbf{X}$ is of full rank, then with $\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}}$ large enough, and with $M_{\pi}\geq c_{\pi}^{-1}$, then with probability converging to $1$, any $\tilde{\gamma}^{*}$ that satisfies $\pi_{i}^{*}\equiv\phi(\tilde{X}_{i}^{\top}\tilde{\gamma}^{*})$ for all $i$ is a feasible solution to (8).
@@
$\hat{\tau}_{\textrm{DCal}}$ is the same as $\hat{\tau}_{\textrm{SCal},r}$ except now $\hat{\pi}$ is also replaced by the post-calibration PS estimator $\tilde{\pi}(x)=\phi(\tilde{\gamma}^{\top}x)$. Constraints (9) to (11) in the program (8) balances $\tilde{\pi}$ such that $T/\tilde{\pi}$ is on average close to $1$, gauged by covariates $\mathbf{X}$ (and the augmented covariates $\mathbf{\tilde{X}}$) and the OR calibrator $\bm{\hat{\mu}}$ under certain data-dependent weights. There is no need to use the auxiliary dataset in the constrained program (8) because the following moment equation holds as long as $\tilde{\pi}_{i}$ does not depend on $Y_{i}$:
@@
Obviously, the program (8) is unrelated to $Y$ from the main dataset. The moment equations (13) and (7) in Remark 1 are direct consequence of the double-robustness of the influence function $\mathsf{IF}(\theta)$ defined in (1). ∎

### v4

@@ 1.2 Notation and organization of the paper
Before moving forwards, we outline some notation that is used throughout. We denote $\theta$ as the nuisance parameters. All expectations and probabilities (e.g. $\mathbb{E}$ and $\mathbb{P}$) are understood to be under the true data generating distribution unless stated otherwise (in that case, we will attach $\theta$ in the subscripts (e.g. $\mathbb{E}_{\theta}$ and $\mathbb{P}_{\theta}$) to emphasize the dependence of the distribution on $\theta$). As will be clear later in Section 2, we may split the entire dataset into three parts, a main dataset $\mathcal{D}$ of size $n$, an auxiliary dataset $\mathcal{D}_{\mathrm{aux}}$ of size $n_{\mathrm{aux}}$, and a training dataset $\mathcal{D}_{\mathrm{tr}}$ of size $n_{\mathrm{tr}}$. For instance, $\mathbf{T}\equiv(T_{1},\cdots,T_{n})^{\top}$ and $\mathbf{X}=(X_{1},\cdots,X_{n})^{\top}$, respectively, denote the treatment variables and the covariates over the main dataset. The quantities computed from a given dataset will be attached, in the subscripts, by the index of that dataset. For example, the $n_{\mathrm{tr}}\times p$ design matrix from the training dataset will be denoted as $\mathbf{X}_{\mathrm{tr}}$, and similarly for $\mathbf{X}_{\mathrm{aux}}$. For each random variable from the training or the auxiliary dataset, we attach a superscript to distinguish them from the main dataset, e.g. $X_{i}^{(\mathrm{aux})}$ is the covariate for the $i$-th subject from the auxiliary dataset. Occasionally, we use the empirical mean operators over the above three datasets, respectively denoted as $\mathbb{P}_{n},\mathbb{P}_{n_{\mathrm{tr}}}$ and $\mathbb{P}_{n_{\mathrm{aux}}}$. We then let $\|X\|_{n,2}\coloneqq(\{\mathbb{P}_{n}X_{1}^{2}\}^{1/2},\cdots,\{\mathbb{P}_{n}X_{p}^{2}\}^{1/2})^{\top}$ denote the column-wise $\ell_{2}$-norms of a given matrix $\mathbf{X}$. With a slight abuse of notation, for an $n\times p$-matrix $\mathbf{V}$ with its rows being $n$ i.i.d. copies $V_{1},\cdots,V_{n}$ of a $p$-dimensional random vector $V$, we let $\|V\|_{n,2}$ denote the $p$-dimensional vector of the column-wise $\ell_{2}$-norms of $\mathbf{V}$. Finally, given two vectors $v_{1},v_{2}$ of the same length, $v_{1}\odot v_{2}$ denotes the component-wise multiplication between $v_{1}$ and $v_{2}$.
In the sequel, we will frequently utilize an $\ell_{\infty}$-calibration approach realized by optimization programs with many inequality constraints. Since different optimization programs will be designed for different parameters, to streamline the notation, we introduce the following $\ell_{\infty}$-calibration operator. To this end, we first define the following two versions of short-hand notation for residuals: given generic random variables $S$ and $g$, $\Delta_{1,S,g}^{(a)}\coloneqq S^{(a)}-g$ and $\Delta_{2,S,g}\coloneqq S/g-1$ where $a\in\{\mathrm{aux},\mathrm{main}\}$ denotes the source of the dataset. As before, if $a=\mathrm{main}$, we silence the superscript. We deliberately choose not to define the second version with superscript $(a)$ because it will not depend on the auxiliary dataset in all appearances in our paper. With these, we define the $\ell_{\infty}$-calibration operator as follows:
$\hat{\bm{\varrho}}\coloneqq\mathcal{L}_{\bm{\varrho},\mathcal{C}_{\bm{\varrho}}}\left(\mathsf{f};\Delta;\mathrm{h};\mathrm{b}\right)\coloneqq{\arg\min}_{\bm{\varrho}\in\mathcal{C}_{\bm{\varrho}}:|\mathbb{P}_{n}[\Delta\cdot\mathrm{h}]|\leq\mathrm{b}}\mathsf{f}(\bm{\varrho}).$ (3)
We now unpack the above definition. Here $\mathcal{C}_{\bm{\varrho}}\subset\mathbb{R}^{k}$ is the constraint set of $\bm{\varrho}$ and $\mathsf{f}:\mathbb{R}^{k}\rightarrow\mathbb{R}_{\geq 0}$ is the objective function. $\Delta$, depending on $\bm{\varrho}$, is the residual that we would like to calibrate such that it satisfies certain constraints that will be clear later. When $\Delta=\Delta_{1,S,g}^{(\mathrm{aux})}$, to avoid clutter, we slightly abuse notation and define $\mathbb{P}_{n}[\Delta_{1,S,g}^{(\mathrm{aux})}\cdot\mathrm{h}]\coloneqq\mathbb{P}_{n_{\mathrm{aux}}}[S^{(\mathrm{aux})}\mathrm{h}^{(\mathrm{aux})}]-\mathbb{P}_{n}[g\mathrm{h}]$. Here $\mathrm{h}\in\mathbb{R}^{l}$ denotes the (random) test functions that the residuals $\Delta^{(a)}$ are calibrated against and $\mathrm{b}\in\mathbb{R}^{l}$ of the same dimension as $\mathrm{h}$ denotes the upper bounds that the calibration aims to fulfill via inequality constraints. As will be clear later, $\mathrm{b}$ can be allowed to be data-dependent, hence possibly random.
@@
The rest of this paper is organized as follows. In Section 2, we illustrate our new methodology for ATE, with the theoretical analysis detailed in Section 3. Section 4 extends our new approach to (1) the regression coefficient estimation in a high-dimensional partially linear model, (2) the estimation of local average treatment effect with instrumental variables with high-dimensional observed confounding variables, and (3) the approximately sparse GLMs considered in Bradic et al. (2019a). In Section 5, for the problem of estimating ATE, we examine the finite-sample performance of a variant of the DCal estimator that is computationally more feasible. Compared against several popular methods, our new method exhibits comparable, and sometimes improved statistical properties and computational runtime. The code for replicating our simulation studies can be accessed via the link https://github.com/Cinbo-Wang/DCal. Lastly, Section 6 concludes this article and discusses some future directions. Some of the technical details are deferred to the Appendix.
@@
Chernozhukov et al. (2022b) $s_{\pi}\wedge s_{r}=o(\sqrt{n}/\log p)$
@@ 2 Average treatment effects: The Double-Calibration estimator
In this section, we illustrate our new “Double-Calibration (DCal)” methodology for ATE estimation. We observe data $\{O_{i}\coloneqq(X_{i},T_{i},Y_{i})\}_{i=1}^{N}$ that are i.i.d. draws from some underlying data-generating distribution $\mathbb{P}_{\theta^{\ast}}$ where from this section onward $\theta^{\ast}\coloneqq(\pi^{\ast},r^{\ast})$ denotes the true OR and PS models. We assume throughout that, as standard in the literature on treatment effect estimation, $\mathbb{P}_{\theta^{\ast}}$ satisfies the unconfoundedness assumption:
@@
where $Y_{i}(1)$ and $Y_{i}(0)$ are two potential outcomes; and the overlap/positivity assumption:
@@
It is well-known that, under Consistency, Conditions 1 and 2, the parameter of interest, ATE, can be identified as $\tau^{\ast}\equiv\mathbb{E}[r^{\ast}(X)]\equiv\mathbb{E}\left[\frac{TY}{\pi^{\ast}(X)}\right]\equiv\mathbb{E}[H(\theta^{\ast})]$. We further impose the following modeling assumption on the nuisance parameters $\theta^{\ast}=(\pi^{\ast},r^{\ast})$:
@@ 2.1 Preliminary: Singly-calibrating the OR model
$\hat{\tau}_{\textrm{SCal},r}\coloneqq\frac{1}{n}\sum_{i=1}^{n}H_{i}(\hat{\pi},\tilde{r})\equiv\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\tilde{r}_{i})}{\hat{\pi}_{i}}+\tilde{r}_{i}\right),\text{ where }\tilde{r}_{i}\coloneqq\hat{r}_{i}+\hat{\mu}_{i}.$ (5)
@@
$\hat{\bm{\mu}}=$ $\;{\arg\min}_{\bm{\mu}\in\mathbb{R}^{n}}\|\bm{\mu}\|_{2}^{2}$ (6a)
$\mathrm{s.t.}\;\;$ $\left\|\frac{1}{n_{\mathrm{aux}}}{\mathbf{X}}_{\mathrm{aux}}^{\top}{\bm{\Pi}}_{\mathrm{aux}}\tilde{\mathbf{R}}_{\mathrm{aux}}-\frac{1}{n}{\mathbf{X}}^{\top}{\bm{\Pi}}{\bm{\mu}}\right\|_{\infty}\leq\eta_{r},$ (6b)
$\left\|{\bm{\mu}}\right\|_{\infty}\leq M_{r}$ (6c)
where $M_{r}>0$ is some known constant that depends on the bounds on $\hat{r}$ and $r^{\ast}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the covariate matrix, ${\bm{\Pi}}\in\mathbb{R}^{n\times n}$ is a diagonal matrix whose $(i,i)$-th entry is $\phi^{\prime}(X_{i}^{\top}\hat{\gamma})/\hat{\pi}_{i}$ and $\tilde{\mathbf{R}}$ is the $n$-dimensional residual vector whose $i$-th element is $\tilde{R}_{i}\coloneqq T_{i}/\hat{\pi}_{i}\cdot(Y_{i}-\hat{r}_{i})$. Recalling from the Notation section, the quantities $\mathbf{X}_{\mathrm{aux}}$ etc. are defined analogously on the auxiliary dataset $\mathcal{D}_{\mathrm{aux}}$. In addition, it is understood that if $\bm{\mu}\equiv\bm{0}$ has already satisfied the constraints (6b) and (6c), the above program will simply output $\hat{\bm{\mu}}=\bm{0}$, or equivalently, $\tilde{r}_{i}\equiv\hat{r}_{i}$ for all $i=1,\cdots,n$.
Since we have introduced the $\ell_{\infty}$-calibration operator (3), we can equivalently represent the above constrained program (6a) to (6c) more compactly as follows:
$\hat{\bm{\mu}}\coloneqq\mathcal{L}_{\bm{\mu},B_{\infty}^{n}(M_{r})}\left(\|\cdot\|_{2}^{2};\Delta^{(\mathrm{aux})}_{1,T(Y-\hat{r})/\hat{\pi},\mu};\phi^{\prime}(X^{\top}\hat{\gamma})X/\hat{\pi};\mathbbm{1}_{p}\eta_{r}\right)$ (7)
where $B_{\infty}^{n}(M_{r})$ denotes the $\ell_{\infty}$-ball in $\mathbb{R}^{n}$ with radius $M_{r}$. The correspondence of the above display to the generic $\ell_{\infty}$-calibration operator $\mathcal{L}_{\bm{\varrho},\mathcal{C}_{\bm{\varrho}}}(\mathsf{f};\Delta;\mathrm{h};\mathrm{b})$ defined previously is as follows:
$\bm{\varrho}$ $\leftarrow\bm{\mu},\mathsf{f}(\cdot)\leftarrow\|\cdot\|_{2}^{2},\mathcal{C}_{\bm{\varrho}}\leftarrow B_{\infty}^{n}(M_{r}),$
$\Delta$ $\leftarrow\Delta^{(\mathrm{aux})}_{1,T(Y-\hat{r})/\hat{\pi},\mu},\mathrm{h}\leftarrow\phi^{\prime}(X^{\top}\hat{\gamma})X/\hat{\pi},\mathrm{b}\leftarrow\mathbbm{1}_{p}\eta_{r}.$
@@
$\hat{\tau}_{\textrm{SCal},r}$ is the same as the DML estimator $\hat{\tau}_{\mathrm{DML}}$ except $\hat{r}$ is replaced by the post-calibration OR estimator $\tilde{r}=\hat{r}+\hat{\mu}$. It is in fact a variant of the recently proposed DIPW estimator for ATE (Wang and Shah, 2024). Constraint (6b) balances $\tilde{r}$ such that it is on average close to $Y$, gauged by covariates $X$ under certain data-dependent weights. The use of the auxiliary dataset $\mathcal{D}_{\mathrm{aux}}$ is to preserve the following moment equations satisfied by the calibrated OR estimator $\tilde{r}$:
@@
because $\hat{\mu}_{i}$’s, as constructed, only depends on $X$, but not on $T$ or $Y$, from the main dataset. Note that (8) in fact holds with any quantity, that is independent of $T_{i}$ given $X_{i}$ from the main dataset $\mathcal{D}$, in place of $\hat{\mu}_{i}$. The above moment equation plays a key role in the analysis of the statistical property of $\hat{\tau}_{\textrm{SCal},r}$ in Wang and Shah (2024). ∎
Using the analysis in the proof of Wang and Shah (2024, Theorem 2), since $\hat{\beta}$ and $\hat{\gamma}$, computed using $\mathcal{D}_{\mathrm{tr}}$, are independent from the main sample $\mathcal{D}$, by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}}$, $\hat{\tau}_{\textrm{SCal},r}$ can be shown to be a $\sqrt{n}$-consistent estimator of $\bar{\tau}^{\ast}$ whenever the PS-sparsity $s_{\pi}\ll\sqrt{n}/\log p$, under Conditions 1 to 3.
@@ 2.2 Double calibration
$\hat{\tau}_{\mathrm{DCal}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}H_{i}(\tilde{\pi},\tilde{r})\equiv\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\tilde{r}_{i})}{\tilde{\pi}_{i}}+\tilde{r}_{i}\right).$
@@
$\tilde{\gamma}\coloneqq$ $\ {\arg\min}_{\gamma}\|\gamma-\hat{\gamma}\|_{1}$ (9a)
@@
$\left|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}}-1\right)\hat{\mu}_{i}\right|\leq\eta_{\pi_{1}}\cdot\frac{\|\hat{\bm{\mu}}\|_{2}}{\sqrt{n}}$ (9c)
@@
$\max_{i}T_{i}/\pi_{i}\leq M_{\pi}.$ (9e)
In this program, $\pi_{i}\equiv\pi(X_{i})\coloneqq\phi(\tilde{X}_{i}^{\top}\gamma)$ and $M_{\pi}>0$ is some known constant that depends on $c_{\pi}$ [^6] In (9b), we use the convention that a function $g:\mathbb{R}^{p}\mapsto\mathbb{R}$ applied to a matrix ${\bm{M}}\in\mathbb{R}^{n\times p}$ ($p$ could be equal to $1$) is a $n$-dimensional vector formed by applying $g$ to each row of the matrix.. Notice also that in the above program, when $n>p$, then with a slight abuse of notation, we let $\hat{\gamma}$ denote the $n$ dimensional vector with first $p$ element equal to the original estimate $\hat{\gamma}$, and the rest $n-p$ coordinates equal to zero. Similar to the OR calibration step, we can compactly represent the above PS calibration step using the $\ell_{\infty}$-calibration operator:
$\tilde{\gamma}\coloneqq\mathcal{L}_{\gamma,\tilde{B}_{\infty}^{p}(M_{\pi})}\left(\|\cdot-\hat{\gamma}\|_{1};\Delta_{2,T,\phi(X^{\top}\gamma)};\mathrm{h};\mathrm{b}\right)$ (10)
where
$\tilde{B}_{\infty}^{p}(M_{\pi})$$\coloneqq\left\{\gamma:\|T/\phi(\tilde{X}^{\top}\gamma)\|_{\infty}\leq M_{\pi}\right\},\mathrm{h}\coloneqq\left(\psi^{\prime}(X^{\top}\hat{\beta})X^{\top},\hat{\mu},\tilde{X}^{\top}\right)^{\top},$
and $\text{and }\mathrm{b}$ $\coloneqq\left(\frac{\eta_{\pi_{1}}}{\sqrt{n}}\|\psi^{\prime}(X^{\top}\hat{\beta})X\|_{n,2},\frac{\eta_{\pi_{1}}\|\hat{\bm{\mu}}\|_{2}}{\sqrt{n}},\frac{\eta_{\pi_{2}}}{\sqrt{n}}\|\tilde{X}\|_{n,2}\right)^{\top},$
where we recall that, for an $n\times p$-matrix $\mathbf{V}$ with its rows being $n$ i.i.d. copies $V_{1},\cdots,V_{n}$ of a $p$-dimensional random vector $V$, we let $\|V\|_{n,2}$ denote the $p$-dimensional vector of the column-wise $\ell_{2}$-norms of $\mathbf{V}$. In particular, if the initial estimator $\hat{\gamma}$ has already satisfied all the inequality constraints, (10) will simply output $\hat{\gamma}$ without further solving the optimization problem.
An appealing property of (9a) is that it is always guaranteed to have a feasible solution, even without any assumptions on the sparsity of the PS or the OR model:
@@
Suppose Condition 2 holds, then with $\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}}$ large enough, and with $M_{\pi}\geq c_{\pi}^{-1}$, we have that given any full rank $\mathbf{X}$ and $(\hat{\bm{\mu}},\hat{\gamma},\hat{\beta})$, with probability converging to $1$, any $\tilde{\gamma}^{*}$ that satisfies $\pi_{i}^{*}\equiv\phi(\tilde{X}_{i}^{\top}\tilde{\gamma}^{*})$ for all $i$ is a feasible solution to (9a).
@@
$\hat{\tau}_{\mathrm{DCal}}$ is the same as $\hat{\tau}_{\textrm{SCal},r}$ except now $\hat{\pi}$ is also replaced by the post-calibration PS estimator $\tilde{\pi}(x)=\phi(\tilde{\gamma}^{\top}x)$. Constraints (9b) to (9d) in the program (9a) balances $\tilde{\pi}$ such that $T/\tilde{\pi}$ is on average close to $1$, gauged by covariates $\mathbf{X}$ (and the augmented covariates $\mathbf{\tilde{X}}$) and the OR calibrator $\bm{\hat{\mu}}$ under certain data-dependent weights. There is no need to use the auxiliary dataset in the constrained program (9a) because the following moment equation holds as long as $\tilde{\pi}_{i}$ does not depend on $Y_{i}$:
@@
Obviously, the program (9a) is unrelated to $Y$ from the main dataset. The moment equations (11) and (8) in Remark 1 are direct consequence of the double-robustness of the influence function $\mathsf{IF}(\theta)$ defined in (1).
It is worth noting that $\hat{\tau}_{\textrm{DCal}}$ calibrates OR and PS in order, and this ordering is vital for the two moment equations (8) and (11) to hold. Were the ordering reversed, with $\hat{\bm{\mu}}$ calibrated using $\tilde{\pi}$ which in turn depends on $\mathbf{T}$ in the main sample, it would not have been as straightforward to show that moment equation (8) holds (even approximately). ∎